OpenStudy (anonymous):

describe how to draw a line that passes through the origin and has a slope of 3/5

OpenStudy (anonymous):

Use point slope form which is y-y1=m(x-x1) Plug in your numbers: y-0=3/5(x-0) Simplify: y=3/5x

OpenStudy (anonymous):

wHAT?

OpenStudy (anonymous):

Look up point slope form. y1 and x1 are your points (which are 0 and 0 because those are the points of the origin) and m is your slope then you just plugin.